# A Cat Is A Cat (Not A Dog!): Unraveling Information Mix-ups in Text-to-Image Encoders through Causal Analysis and Embedding Optimization

### v4

@@ Appendix F Image quality evaluation
We follow SD3 [2] to conduct the image quality evaluation on Fréchet Inception Distance (FID) with CLIP L/14 image features on the generated images and the COCO 2017 val dataset [9] in 5,000 samples. The FID of (SD 1.4, SD 1.4 + TEBOpt), (SDXL-Turbo, SDXL-Turbo + TEBOpt), and (SD3, SD3 + TEBOpt) are (133.08, 133.30), (202.50, 200.71), and (143.77, 142.20), where FIDs are higher than we usually see from text-to-image models is because the 5,000 generated sets are based on the plain prompt structure "a <objA> and a <objB>". Visual performance is mainly affected by the selected text-to-image model as the FID for the generated images w/ or w/o TEBOpt are within marginal differences in the same model. When these 3 models work with TEBOpt, only SD 1.4 gets a 0.22 increase in FID score, while SDXL-Turbo and SD3 result in a 1.79 and 1.57 decrease in FID scores. It proves that TEBOpt would improve the general image quality.
@@ Appendix G Interesting questions aroused by reviewers
We explored this question when we hypothesized the solution to our proposed problem. We conducted an experiment on the same 400-prompt set described in the paper, eliminating "all" earlier information from accumulating in subsequent tokens. The results are presented in the table below. Without shared embeddings across tokens, the generation process failed to produce co-existing objects. Specifically, during the denoising process, each object token responded in the central region, as observed in the cross-attention maps, resulting in no object co-existence. In conclusion, maintaining a proper proportion of earlier object token information in the latter tokens (excluding those with concrete meanings) has more positive than negative effects, especially in generating co-existing objects within a given prompt. Therefore, we propose to optimize the critical tokens’ embeddings in the paper.
@@ Appendix H The instruction for participants in human evaluation for our proposed evaluation metric
"You will be given an image and a text description. The text is described as "a/an <object1> and a/an <object2>". Please determine whether the objects in the image are a combination of object 1 and object 2, or if any object is missing. Select the corresponding answer from the 5 options: i) two objects exist, ii) mixture exists, iii) missing object 1, iv) missing object 2, or v) no objects exist."

### v5

@@ Appendix F Image quality evaluation
We follow SD3 [2] to conduct the image quality evaluation on Fréchet Inception Distance (FID) with CLIP L/14 image features on the generated images and the COCO 2017 val dataset [8] in 5,000 samples. The FID of (SD 1.4, SD 1.4 + TEBOpt), (SDXL-Turbo, SDXL-Turbo + TEBOpt), and (SD3, SD3 + TEBOpt) are (133.08, 133.30), (202.50, 200.71), and (143.77, 142.20), where FIDs are higher than we usually see from text-to-image models is because the 5,000 generated sets are based on the plain prompt structure "a <objA> and a <objB>". This experiment proves that visual performance is mainly affected by the selected text-to-image model as the FID for the generated images w/ or w/o TEBOpt are within marginal differences in the same model. When these 3 models work with TEBOpt, only SD 1.4 gets a 0.22 increase in FID score, while SDXL-Turbo and SD3 result in a 1.79 and 1.57 decrease in FID scores. It proves that TEBOpt would improve the general image quality.
@@ Appendix G Interesting questions aroused by reviewers
We explored this question when we hypothesized the solution to our proposed problem. We conducted an experiment on the same 400-prompt set described in the paper, eliminating "all" earlier information from accumulating in subsequent tokens. The results are presented in Table 9. Without shared embeddings across tokens, the generation process failed to produce co-existing objects. Specifically, during the denoising process, each object token responded in the central region, as observed in the cross-attention maps, resulting in no object co-existence. In conclusion, maintaining a proper proportion of earlier object token information in the latter tokens (excluding those with concrete meanings) has more positive than negative effects, especially in generating co-existing objects within a given prompt. Therefore, we propose to optimize the critical tokens’ embeddings in the paper.
@@
Info bias 1.18
Table 9: Evaluation for eliminating all information accumulation.
@@ Appendix H The instruction for participants in human evaluation for our proposed evaluation metric
"You will be given an image and a text description. The text is described as "a/an <object1> and a/an <object2>". Please determine whether the objects in the image are a combination of object 1 and object 2, or if any object is missing. Select the corresponding answer from the 5 options: i) two objects exist, ii) mixture exists, iii) missing object 1, iv) missing object 2, or v) no objects exist." The screenshot of the human evaluation is demonstrated in Fig. 15.